# CBF-RL: Safety Filtering Reinforcement Learning in Training with Control Barrier Functions

### v2

@@ Abstract
Reinforcement learning (RL), while powerful and expressive, can often prioritize performance at the expense of safety. Yet safety violations can lead to catastrophic outcomes in real-world deployments. Control Barrier Functions (CBFs) offer a principled method to enforce dynamic safety—traditionally deployed online via safety filters. While the result is safe behavior, the fact that the RL policy does not have knowledge of the CBF can lead to conservative behaviors. This paper proposes CBF-RL, a framework for generating safe behaviors with RL by enforcing CBFs in training. CBF-RL has two key attributes: (1) minimally modifying a nominal RL policy to encode safety constraints via a CBF term, (2) and safety filtering of the policy rollouts in training. Theoretically, we prove that continuous-time safety filters can be deployed via closed-form expressions on discrete-time roll-outs. Practically, we demonstrate that CBF-RL internalizes the safety constraints in the learned policy—both enforcing safer actions and biasing towards safer rewards—enabling safe deployment without the need for an online safety filter. We validate our framework through ablation studies on navigation tasks and on the Unitree G1 humanoid robot, where CBF-RL enables safer exploration, faster convergence, and robust performance under uncertainty, enabling the humanoid robot to avoid obstacles and climb stairs safely in real-world settings without a runtime safety filter.
@@ I Introduction
Humanoid robots are capable of interacting with environments designed for humans. However, the complex environment, high-dimensional robot dynamics, and noise of the sensors also make them highly vulnerable to unsafe control inputs. One unsafe action could lead to damage to both the robot and its surroundings, and thus ensuring safety is essential. Meanwhile, reinforcement learning (RL) has emerged as a powerful tool for humanoid robots to achieve diverse skills, but focuses mostly on performance [1, 2, 3] and expressiveness [4, 5, 6, 7].
@@
In this paper, we propose integrating formal safety mechanisms with the powerful exploration and exploitation abilities of RL so that learned policies can reduce or prevent catastrophic behaviors. To achieve this, we turn to Control Barrier Functions (CBFs) [8] for a principled way to encode state-based safety constraints as forward-invariant sets. The CBF conditions are often enforced using safety filters [9]: quadratic programs satisfying the safety constraint by minimally modifying a proposed control input.
There are two key approaches to instantiating safety filters in RL. The first approach is safety filtering the RL-proposed action and projects it into the safe set before execution [10, 11, 12, 13] or performing constrained updates of the gradient [14, 15]. This guarantees safety at runtime, but the filter must remain in the loop at deployment, and the learned policy may never internalize the constraint. This prevents a high-dimensional agent, like a humanoid robot, from discovering novel or efficient behaviors since the exploration space is pruned too aggressively. Both also require solving an optimization program at every control step, which may be computationally expensive. The second approach is reward shaping where a residual augments the reward term to penalize states that approach or violate constraint boundaries [16, 17, 18, 19, 20, 21] and encourages the agent towards safer behaviors without active filtering. This alone does not directly enforce safe actions during training and is often sensitive to the choice of penalty weights, possibly being insufficient in safety-critical applications. This paper proposes a fusion of these two approaches to integrating CBFs into RL.
@@ I-A Contributions
We evaluate CBF-RL, and its dual approach to safe RL, with ablations using a 2D navigation task with dynamics randomness, analyzing task completion rates and robustness tests. We also train humanoid locomotion policies in IsaacLab [22] with full-order dynamics and domain randomization for obstacle avoidance and stair climbing tasks. The approach is validated on hardware using a Unitree G1 humanoid with zero-shot sim-to-real policies to exhibit the effectiveness of the proposed framework in high-dimensional, complex systems. With the dual-trained policy, the robot can successfully navigate around obstacles and climb stairs under command sequences that would lead to failure with nominal policies that don’t leverage the CBF-RL framework.
@@ I-B Related Work
Due to the ease of modifying rewards for training, a number of works incorporate safety value functions into the reward structure to guide policies toward safety. [16, 17] use a fixed penalty, [18, 19, 21] utilize correction-proportional penalties, and [20] proposes an explicit barrier-inspired reward shaping mechanism to reduce unsafe exploration. These methods encourage safe behavior but do not explicitly direct the policy to exhibit safe behaviors during learning, instead solely relying on the policy to discover safer actions on its own, leading to slower training.
@@
To address this, runtime CBF-based safety filters during training minimally modify the actions of the policy such that the system stays in the user-defined forward-invariant safe set typically by solving an optimization program, be it discrete-time due to the nature of RL training [10, 11, 15, 12, 23, 13], or continuous-time [24, 25, 26] which empirically shows improved safety. For humanoid locomotion, these methods are not ideal as humanoid robots have tight real-time and computation power constraints and inaccurate state estimation from sensor noise. We instead filter only in simulation during training, and show that the resulting policy retains safety even without a runtime filter. We further provide theoretical verification that under certain conditions, continuous-time CBF can be used as conditions for forward invariance for RL simulations that are discrete in nature, and thus we can use the closed-form solution to the CBF-QP to accelerate each step in training.
@@
Many papers utilize model-based approaches [11, 15, 24, 27]. which require access to accurate dynamics models. While theoretically appealing, they are less practical for high-dimensional humanoids where dynamics are complex and uncertain. Some works also do constrained updates of the gradient to ensure that the policy remains safe [14, 15]; however, they also require solving optimization programs at each gradient update step and limit the exploration of the policy. Our framework is model-free, requiring only derivatives of the reduced-order model (e.g. for the kinematics of a humanoid robot, its Jacobian $J$), and emphasizes lightweight integration with standard policy-gradient RL, in our case proximal policy optimization(PPO) [28]. This also leads to the benefit of the policy being able to venture closer to the constraint boundaries, ensuring rich exploration.
@@
In response to the issue of stochastic systems, robust extensions address uncertainty in dynamics or sensing through disturbance observers, Gaussian Process models, or robustified CBFs [29, 25, 24, 26]. These methods improve reliability under uncertainty but add significant complexity and computational burden. Here we show that by relying on domain randomization during training, dual-trained policies remain safer under uncertainty of the system without explicit models.
@@
Another line of work learns barrier-like certificates or relates value functions to barrier properties [30, 31]. These methods aim to automate the design of barrier functions or embed them in differentiable layers. Our approach assumes analytic barrier functions and focuses on pragmatic integration with RL training rather than barrier discovery.
@@
The above works have been applied to domains such as spacecraft inspection [12], drone control [13], autonomous driving [23], and driver assistance [24]. Much of the prior work captures part of the safety challenge, but none directly combines filtering and shaping in a way that allows a policy to internalize safety during training and then act autonomously without a filter at deployment, especially not on a high-dimensional system such as a humanoid robot. This distinction defines the novelty of our contribution.
@@ II Background
Reinforcement Learning We consider the standard RL formulation of a Markov decision process (MDP) $(\mathcal{X},\mathcal{U},P,r,\gamma)$. At each timestep $k$, an agent selects an action $\mathbf{u_{k}}\in\mathcal{U}$, according to a policy $\mathbf{\pi_{\theta}}(\mathbf{u_{k}}|\mathbf{x_{k}})$ based on an observed state $\mathbf{x_{k}}\in\mathcal{X}$, and receives a reward $r(\mathbf{x_{k}},\mathbf{u_{k}})$. The environment follows the transition dynamics $\mathbf{x_{k+1}}\sim P(\cdot|\mathbf{x_{k}},\mathbf{u_{k}})$. The goal is then to maximize the expected discounted return $\mathbb{E}\!\left[\sum_{k=0}^{\infty}\gamma^{t}r(\mathbf{x_{k}},\mathbf{u_{k}})\right]$. During deployment, actions are selected as the conditional expectation $\mathbf{u_{k}}=E[\mathbf{\pi_{\theta}}(\mathbf{u_{k}}|\mathbf{x_{k}})]$. In this work, we specifically use model-free policy-gradient actor-critic methods such as PPO [28], though our approach is agnostic to the specific RL algorithm. RL also often suffers from reward sparsity when unsafe events like obstacle collisions are rare. Thus the policy seldom experiences consequences of unsafe actions, leading to vanishing gradients and unstable, slow training. As such, one part of our method also does reward densification to mitigate this by designing $r(\mathbf{x_{k}},\mathbf{u_{k}})$ to give informative nonterminal signals related to safety.
@@
Reduced-Order Models Consider a continuous-time system with dynamics $\dot{\mathbf{x}}=\mathbf{\phi}(\mathbf{x},\mathbf{u}),$ where $\mathbf{x}\in\mathbb{R}^{n}$ is the state and $\mathbf{u}\in\mathbb{R}^{m}$ is the control input. In our case the state $\mathbf{x}$ may be very high-dimensional, e.g. joint positions and velocities of a robot. Thus we consider a reduced-order state $\mathbf{q}\in\mathbb{R}^{n_{q}}$, $n_{q}<n$, representing key lower-dimensional features such as the robot’s center of mass position. We define a projection $\mathbf{p}:\mathbb{R}^{n}\to\mathbb{R}^{n_{q}}$ that projects the full-order state onto the reduced-order state. Given a locally Lipschitz continuous feedback control law $\mathbf{v}=\mathbf{k}(\mathbf{q})$, the reduced state $\mathbf{q}$ is governed by
@@
where $\mathbf{\psi}(\mathbf{x},\mathbf{v})$ is a control interface lifting the reduced-order input $\mathbf{v}$ to a full-order input $\mathbf{u}$. See [30, 32] for the connections between reduced and full order models.
Control Barrier Function Safety Filters In the control barrier function framework, a set of “safe states” for the system, $\mathcal{S}\subseteq\mathbb{R}^{n}$, is encoded as the zero superlevel set of a continuously differentiable function $h:\mathbb{R}^{n}\to\mathbb{R}$,
$\mathcal{S}\coloneqq\left\{\mathbf{q}\in\mathbb{R}^{n}\,\big|\,h(\mathbf{q})\geq 0\right\},$ (2)
$\partial\mathcal{S}\coloneqq\left\{\mathbf{q}\in\mathbb{R}^{n}\,\big|\,h(\mathbf{q})=0\right\},$ (3)
$\mathrm{int}\left(\mathcal{S}\right)\coloneqq\left\{\mathbf{q}\in\mathbb{R}^{n}\,\big|\,h(\mathbf{q})>0\right\}.$ (4)
@@
A set $\mathcal{S}\subseteq\mathbb{R}^{n}$ is forward invariant for (1) if, for every initial state $\mathbf{q}(t_{0})\in\mathcal{S}$, the resulting state trajectory $\mathbf{q}:I\subseteq\mathbb{R}\to\mathbb{R}^{n}$ remains in $\mathcal{S}$ for all $t\in I\cap\mathbb{R}_{\geq t_{0}}$.
For control-affine systems as in (1), the forward invariance of $\mathcal{S}$ can be enforced using CBFs [8, 9].
@@
Let $\mathcal{S}\subset\mathbb{R}^{n}$ be a set as in (2), with $h$ satisfying $\nabla h|_{\partial\mathcal{S}}\neq 0$. Then, $h$ is a control barrier function (CBF) on $\mathbb{R}^{n}$ if there exists [^1] A function $\alpha:\mathbb{R}\to\mathbb{R}$ belongs to $\mathcal{K}_{\infty}^{e}$ if it is increasing, continuous, and satisfies $\alpha(0)=0,\;\lim_{r\to\pm\infty}\alpha(r)=\pm\infty$. a $\gamma\in\mathcal{K}_{\infty}^{e}$ such that for all $\mathbf{q}\in\mathbb{R}^{n}$,
@@
Any locally Lipschitz controller $\mathbf{k}$ for (1) for which $\mathbf{k}(\mathbf{q})\in\mathcal{U}_{\mathrm{CBF}}(\mathbf{q})\;\forall\mathbf{q}\in\mathcal{S}$ enforces forward invariance of $\mathcal{S}$ [8].
@@
where $\mathbf{F}:\mathbb{R}^{n}\times\mathbb{R}^{m}\to\mathbb{R}^{m}$ is the discretization of(1) over a time interval $\mathrm{\Delta t}>0$ for a constant input $\mathbf{v}$. Here, we focus on enforcing forward invariance at sample times [^2] See [33] for a discussion of zero-order-hold, intersample safety. $k\mathrm{\Delta t}$. This can be achieved using a discrete-time CBF [34, 35].
@@
Let $\mathcal{S}\subseteq\mathbb{R}^{n}$ be as in (2) and $\rho\in[0,1]$. The function $h$ is a discrete-time control barrier function (DTCBF) for (7) if $\forall\mathbf{q}\in\mathcal{S}$ there exists a $\mathbf{v}\in\mathbb{R}^{m}$ for which
@@
Similar to CBFs and continuous-time systems, DTCBFs keep the discrete-time system (7) safe for each $k\in\mathbb{Z}_{\geq 0}$. Here the value of $h$ is lower bounded by a geometrically decaying curve, $h(\mathbf{q}_{k})\geq\rho^{k}h(\mathbf{q}_{0})$ [34]. Thus, they can be used to generate safe control actions through an optimization program wherein a desired but potentially unsafe input $\mathbf{v}^{\textup{des}}_{k}\in\mathbb{R}^{m}$ is minimally modified to produce a safe input:

### v4

@@ Abstract
Reinforcement learning (RL), while powerful and expressive, can often prioritize performance at the expense of safety. Yet safety violations can lead to catastrophic outcomes in real-world deployments. Control Barrier Functions (CBFs) offer a principled method to enforce dynamic safety—traditionally deployed online via safety filters. While the result is safe behavior, the fact that the RL policy does not have knowledge of the CBF can lead to conservative behaviors. This paper proposes CBF-RL, a framework for generating safe behaviors with RL by enforcing CBFs in training. CBF-RL has two key attributes: (1) minimally modifying a nominal RL policy to encode safety constraints via a CBF term, and (2) safety filtering of the policy rollouts in training. Theoretically, we prove that continuous-time safety filters can be deployed via closed-form expressions on discrete-time roll-outs. Practically, we demonstrate that CBF-RL internalizes the safety constraints in the learned policy—both enforcing safer actions and biasing towards safer rewards—enabling safe deployment without the need for an online safety filter. We validate our framework through ablation studies on navigation tasks and on the Unitree G1 humanoid robot, where CBF-RL enables safer exploration, faster convergence, and robust performance under uncertainty, enabling the humanoid robot to avoid obstacles and climb stairs safely in real-world settings without a runtime safety filter.
@@ I Introduction
Humanoid robots are capable of interacting with environments designed for humans. However, the complex environment, high-dimensional robot dynamics, and noise of the sensors also make them highly vulnerable to unsafe control inputs. One unsafe action could lead to damage to both the robot and its surroundings, and thus ensuring safety is essential. Meanwhile, reinforcement learning (RL) has emerged as a powerful tool for humanoid robots to achieve diverse skills, but focuses mostly on performance [crowley2025optimizing, li2025reinforcement, peng2025gait] and expressiveness [he2025asap, allshire2025visual, truong2025beyondmimic, su2025hitter].
@@
In this paper, we propose integrating formal safety mechanisms with the powerful exploration and exploitation abilities of RL so that learned policies can reduce or prevent catastrophic behaviors. To achieve this, we turn to Control Barrier Functions (CBFs) [ames2016control] for a principled way to encode state-based safety constraints as forward-invariant sets. The CBF conditions are often enforced using safety filters [ames2019control]: quadratic programs satisfying the safety constraint by minimally modifying a proposed control input.
There are two key approaches to instantiating safety filters in RL. The first approach is safety filtering the RL-proposed action and projects it into the safe set before execution [cheng_end--end_2019, ma_model-based_2021, van_wijk_safe_2024, bejarano_safety_2025] or performing constrained updates of the gradient [alshiekh_safe_2017, zhang_model-based_2021]. This guarantees safety at runtime, but the filter must remain in the loop at deployment, and the learned policy may never internalize the constraint. This prevents a high-dimensional agent, like a humanoid robot, from discovering novel or efficient behaviors since the exploration space is pruned too aggressively. Both also require solving an optimization program at every control step, which may be computationally expensive. The second approach is reward shaping where a residual augments the reward term to penalize states that approach or violate constraint boundaries [krasowski2022provably, dunlap2023run, wabersich2021predictive, wang2022ensuring, nilaksh_barrier_2024, wang2025omni] and encourages the agent towards safer behaviors without active filtering. This alone does not directly enforce safe actions during training and is often sensitive to the choice of penalty weights, possibly being insufficient in safety-critical applications. This paper proposes a fusion of these two approaches to integrating CBFs into RL.
@@ I-A Contributions
We evaluate CBF-RL, and its dual approach to safe RL, with ablations using a 2D navigation task with dynamics randomness, analyzing task completion rates and robustness tests. We also train humanoid locomotion policies in IsaacLab [mittal2023orbit] with full-order dynamics and domain randomization for obstacle avoidance and stair climbing tasks. The approach is validated on hardware using a Unitree G1 humanoid with zero-shot sim-to-real policies to exhibit the effectiveness of the proposed framework in high-dimensional, complex systems. With the dual-trained policy, the robot can successfully navigate around obstacles and climb stairs under command sequences that would lead to failure with nominal policies that don’t leverage the CBF-RL framework.
@@ I-B Related Work
Due to the ease of modifying rewards for training, a number of works incorporate safety value functions into the reward structure to guide policies toward safety. [krasowski2022provably, dunlap2023run] use a fixed penalty, [wabersich2021predictive, wang2022ensuring, wang2025omni] utilize correction-proportional penalties, and [nilaksh_barrier_2024] proposes an explicit barrier-inspired reward shaping mechanism to reduce unsafe exploration. These methods encourage safe behavior but do not explicitly direct the policy to exhibit safe behaviors during learning, instead solely relying on the policy to discover safer actions on its own, leading to slower training.
@@
To address this, runtime CBF-based safety filters during training minimally modify the actions of the policy such that the system stays in the user-defined forward-invariant safe set typically by solving an optimization program, be it discrete-time due to the nature of RL training [cheng_end--end_2019, ma_model-based_2021, zhang_model-based_2021, van_wijk_safe_2024, zhang_control_2025, bejarano_safety_2025], or continuous-time [hailemichael_safe_2022, emam_safe_2022, cheng_safe_2023] which empirically shows improved safety. For humanoid locomotion, these methods are not ideal as humanoid robots have tight real-time and computation power constraints and inaccurate state estimation from sensor noise. We instead filter only in simulation during training, and show that the resulting policy retains safety even without a runtime filter. We further provide theoretical verification that under certain conditions, continuous-time CBF can be used as conditions for forward invariance for RL simulations that are discrete in nature, and thus we can use the closed-form solution to the CBF-QP to accelerate each step in training.
@@
Many papers utilize model-based approaches [ma_model-based_2021, zhang_model-based_2021, hailemichael_safe_2022, du_reinforcement_2023]. which require access to accurate dynamics models. While theoretically appealing, they are less practical for high-dimensional humanoids where dynamics are complex and uncertain. Some works also do constrained updates of the gradient to ensure that the policy remains safe [alshiekh_safe_2017, zhang_model-based_2021]; however, they also require solving optimization programs at each gradient update step and limit the exploration of the policy. Our framework is model-free, requiring only derivatives of the reduced-order model (e.g. for the kinematics of a humanoid robot, its Jacobian $J$), and emphasizes lightweight integration with standard policy-gradient RL, in our case proximal policy optimization(PPO) [schulman2017proximal]. This also leads to the benefit of the policy being able to venture closer to the constraint boundaries, ensuring rich exploration.
@@
In response to the issue of stochastic systems, robust extensions address uncertainty in dynamics or sensing through disturbance observers, Gaussian Process models, or robustified CBFs [li_robust_2020, emam_safe_2022, hailemichael_safe_2022, cheng_safe_2023]. These methods improve reliability under uncertainty but add significant complexity and computational burden. Here we show that by relying on domain randomization during training, dual-trained policies remain safer under uncertainty of the system without explicit models.
@@
Another line of work learns barrier-like certificates or relates value functions to barrier properties [cohen_safe_2023, tan_value_2023]. These methods aim to automate the design of barrier functions or embed them in differentiable layers. Our approach assumes analytic barrier functions and focuses on pragmatic integration with RL training rather than barrier discovery.
@@
The above works have been applied to domains such as spacecraft inspection [van_wijk_safe_2024], drone control [bejarano_safety_2025], autonomous driving [zhang_control_2025], and driver assistance [hailemichael_safe_2022]. Much of the prior work captures part of the safety challenge, but none directly combines filtering and shaping in a way that allows a policy to internalize safety during training and then act autonomously without a filter at deployment, especially not on a high-dimensional system such as a humanoid robot. This distinction defines the novelty of our contribution.
@@ II Background
Reinforcement Learning We consider the standard RL formulation of a Markov decision process (MDP) $(\mathcal{X},\mathcal{U},P,r,\gamma)$. At each timestep $k$, an agent selects an action $\mathbf{u_{k}}\in\mathcal{U}$, according to a policy $\mathbf{\pi_{\theta}}(\mathbf{u_{k}}|\mathbf{x_{k}})$ based on an observed state $\mathbf{x_{k}}\in\mathcal{X}$, and receives a reward $r(\mathbf{x_{k}},\mathbf{u_{k}})$. The environment follows the transition dynamics $\mathbf{x_{k+1}}\sim P(\cdot|\mathbf{x_{k}},\mathbf{u_{k}})$. The goal is then to maximize the expected discounted return $\mathbb{E}\!\left[\sum_{k=0}^{\infty}\gamma^{t}r(\mathbf{x_{k}},\mathbf{u_{k}})\right]$. During deployment, actions are selected as the conditional expectation $\mathbf{u_{k}}=E[\mathbf{\pi_{\theta}}(\mathbf{u_{k}}|\mathbf{x_{k}})]$. In this work, we specifically use model-free policy-gradient actor-critic methods such as PPO [schulman2017proximal], though our approach is agnostic to the specific RL algorithm. RL also often suffers from reward sparsity when unsafe events like obstacle collisions are rare. Thus the policy seldom experiences consequences of unsafe actions, leading to vanishing gradients and unstable, slow training. As such, one part of our method also does reward densification to mitigate this by designing $r(\mathbf{x_{k}},\mathbf{u_{k}})$ to give informative nonterminal signals related to safety.
@@
Reduced-Order Models Consider a continuous-time system with dynamics $\dot{\mathbf{x}}=\mathbf{\phi}(\mathbf{x},\mathbf{u}),$ where $\mathbf{x}\in\mathbb{R}^{n}$ is the state and $\mathbf{u}\in\mathbb{R}^{m}$ is the control input. In our case the state $\mathbf{x}$ may be very high-dimensional, e.g. joint positions and velocities of a robot. Thus we consider a reduced-order state $\mathbf{q}\in\mathbb{R}^{n_{q}}$, $n_{q}<n$, representing key lower-dimensional features such as the robot’s center of mass position. We define a projection $\mathbf{p}:\mathbb{R}^{n}\to\mathbb{R}^{n_{q}}$ that projects the full-order state onto the reduced-order state. Given a locally Lipschitz continuous feedback control law $\mathbf{v}=\mathbf{k}(\mathbf{q})$, the reduced-order state $\mathbf{q}$ is
@@
where $\mathbf{\psi}(\mathbf{x},\mathbf{v})$ is a control interface lifting the reduced-order input $\mathbf{v}$ to a full-order input $\mathbf{u}$. See [cohen_safe_2023, cohen2025safety] for the connections between reduced and full order models.
Control Barrier Function Safety Filters In the control barrier function framework, a set of “safe states” for the system, $\mathcal{S}\subseteq\mathbb{R}^{n_{q}}$, is encoded as the zero superlevel set of a continuously differentiable function $h:\mathbb{R}^{n_{q}}\to\mathbb{R}$,
$\mathcal{S}\coloneqq\left\{\mathbf{q}\in\mathbb{R}^{n_{q}}\,\big|\,h(\mathbf{q})\geq 0\right\},$ (2)
$\partial\mathcal{S}\coloneqq\left\{\mathbf{q}\in\mathbb{R}^{n_{q}}\,\big|\,h(\mathbf{q})=0\right\},$ (3)
$\mathrm{int}\left(\mathcal{S}\right)\coloneqq\left\{\mathbf{q}\in\mathbb{R}^{n_{q}}\,\big|\,h(\mathbf{q})>0\right\}.$ (4)
@@
A set $\mathcal{S}\subseteq\mathbb{R}^{n_{q}}$ is forward invariant for (1) if, for every initial state $\mathbf{q}(t_{0})\in\mathcal{S}$, the resulting state trajectory $\mathbf{q}:I\subseteq\mathbb{R}\to\mathbb{R}^{n_{q}}$ remains in $\mathcal{S}$ for all $t\in I\cap\mathbb{R}_{\geq t_{0}}$.
For control-affine systems as in (1), the forward invariance of $\mathcal{S}$ can be enforced using CBFs [ames2016control, ames2019control].
@@
Let $\mathcal{S}\subset\mathbb{R}^{n_{q}}$ be a set as in (2), with $h$ satisfying $\nabla h|_{\partial\mathcal{S}}\neq 0$. Then, $h$ is a control barrier function on $\mathbb{R}^{n_{q}}$ if there exists1111: A function $\alpha:\mathbb{R}\to\mathbb{R}$ belongs to $\mathcal{K}_{\infty}^{e}$ if it is increasing, continuous, and satisfies $\alpha(0)=0,\;\lim_{r\to\pm\infty}\alpha(r)=\pm\infty$.a $\gamma\in\mathcal{K}_{\infty}^{e}$ such that for all $\mathbf{q}\in\mathbb{R}^{n_{q}}$,
@@
Any locally Lipschitz controller $\mathbf{k}$ for (1) for which $\mathbf{k}(\mathbf{q})\in\mathcal{U}_{\mathrm{CBF}}(\mathbf{q})\;\forall\mathbf{q}\in\mathcal{S}$ enforces forward invariance of $\mathcal{S}$ [ames2016control].
@@
where $\mathbf{F}:\mathbb{R}^{n_{q}}\to\mathbb{R}^{n_{q}}$ and $\mathbf{G}:\mathbb{R}^{n_{q}}\to\mathbb{R}^{n_{q}\times m}$ are the discretization of (1) over a time interval $\mathrm{\Delta t}>0$ for a constant input $\mathbf{v}$. Here, we focus on enforcing forward invariance at sample times2222: See [breeden2021control] for a discussion of zero-order-hold, intersample safety. $k\mathrm{\Delta t}$. This can be achieved using a discrete-time CBF [agrawal2017discrete, ahmadi2019safe].
@@
Let $\mathcal{S}\subseteq\mathbb{R}^{n_{q}}$ be as in (2) and $\rho\in[0,1]$. The function $h$ is a discrete-time control barrier function (DTCBF) for (7) if $\forall\mathbf{q}\in\mathcal{S}$ there exists a $\mathbf{v}\in\mathbb{R}^{m}$ for which
@@
Similar to CBFs and continuous-time systems, DTCBFs keep the discrete-time system (7) safe for each $k\in\mathbb{Z}_{\geq 0}$. Here the value of $h$ is lower bounded by a geometrically decaying curve, $h(\mathbf{q}_{k})\geq\rho^{k}h(\mathbf{q}_{0})$ [agrawal2017discrete]. Thus, they can be used to generate safe control actions through an optimization program wherein a desired but potentially unsafe input $\mathbf{v}^{\textup{des}}_{k}\in\mathbb{R}^{m}$ is minimally modified to produce a safe input: